OpenStudy (anonymous):

-7/b 6/3b^2 Least common denominator

OpenStudy (anonymous):

\(3b^3\)

OpenStudy (anonymous):

actually if it is \(\frac{6}{3b^2}=\frac{2}{b^2}\) you can use \(b^2\) as a least common denominator